OpenStudy (anonymous):

Let y = 20,000(0.95)x represent the buying power of $20,000 with an inflation rate of 5% per year. The table below represents the first 4 years later. Which year is incorrect? Options: A. 1 B. 2 C. 3 D. 4

OpenStudy (anonymous):

here it is!! (:

OpenStudy (anonymous):

please do not criticize (:

OpenStudy (anonymous):

they shouldnt have given you the formula, they just should have given the inflation rate, to make a tiny bit harder also why did u give me a medal on the other question :P

OpenStudy (anonymous):

idk >.< i i just did

OpenStudy (anonymous):

guess i shouldnt give u one on this one then huh?

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

its quite simple check : 1. does 20000 x (0.95) = 19100 2. does 20000 x (0.95)^2 = 18050 3. does 20000 x (0.95)^3 = 17148 4. does 20000 x (0.95)^4 = 16290

OpenStudy (anonymous):

its simple calculator work

OpenStudy (anonymous):

whichever one isnt true is your answer

OpenStudy (anonymous):

Year 1 is incorrect not that hard :|

OpenStudy (anonymous):

20000 x 0.95 =19000 , NOT 19100

OpenStudy (anonymous):

How old are u? Either your a college person or your judgmental. either one im still thankful (:

OpenStudy (anonymous):

im 19 and a half , but isnt this like the easiest question you will ever come across

OpenStudy (anonymous):

its just calculator work

OpenStudy (anonymous):

ik >.< but i cant focus and im barely keeping my eyes open so bear with me pls